OpenStudy (anonymous):

int [1, 4] dx/(x-2)^2 ?

OpenStudy (shubhamsrg):

Substitute (x-2) = t

OpenStudy (anonymous):

\[\int\limits_{1}^{4} \frac{ dx }{ (x-2)^{2} }\]

OpenStudy (anonymous):

it should be infinite

OpenStudy (amoodarya):

its not integrable on [1,4] because of "the function has vertica retricemptote in this interval "

OpenStudy (anonymous):

yes, how do i show that?

OpenStudy (amoodarya):

but in complex you can calcute it !

OpenStudy (shubhamsrg):

Oh yes!

OpenStudy (amoodarya):